OpenStudy (anonymous):

The sum of two number is 21. Twice the larger number is four times the smaller number. Find both numbers. .. Ok so I understand A+B=21 is the first expression and I think the second expression reads as 2b=4a. Now what?

OpenStudy (***[isuru]***):

Hi, u have almost finished the work bro... 2B = 4A now get an expression for either A or B ... let's take A here A = 2B/4 =B/2 now substitute B/2 = A in A + B = 21 and u could get the answer with ur eyes shut!!!! hope this will help ya !

OpenStudy (***[isuru]***):

Did u get it or do u need help in simplification ..:-) ?

OpenStudy (anonymous):

I'm rewriting it now to make sure I understand. My apologies, it takes me a bit to process information-results of a bad car accident, so I'm a bit slower that most

OpenStudy (***[isuru]***):

Oh ! I very sorry to hear that ... u can take ur time

OpenStudy (anonymous):

oh that's ok!! - I think I may have done this backwards... lol I have it reading as 2(42)=4(42/2) which makes it 84 on both sides.

OpenStudy (***[isuru]***):

Let me give u a hand ... A + B = 21 we found out that A = (B/2) so B/2 + B = 21 now solve this expression for B

OpenStudy (anonymous):

14?

OpenStudy (***[isuru]***):

u got it bro... u r correct

OpenStudy (anonymous):

14/2 +b=21 reduced to 7 . subtract 7 from each side B=14

OpenStudy (***[isuru]***):

yep!